OpenStudy (anonymous):

Two years ago a father was four times as old as his son three years from now he will be only three times as old as his son how old is each now?

OpenStudy (anonymous):

Let present age of father be F and present age of son be S. As per the problem, (F-2) = 4*(S-2) and (F+3) = 3*(S+3) Solving for F and S, F = 42 S = 12